OpenStudy (anonymous):

Write down the intensity equation for a two-slit interference pattern with diffraction effects included. Indicate which part originates from double-slit interference and which part from single-slit diffraction. What happens when the angle = 0

OpenStudy (anonymous):

I was given the following \[I _{Double}=4Icos^2(\frac{ \pi*d }{ \lambda*L })y\] \[I_{single}=I _{0}\frac{ \sin^2\frac{ \pi*a*\sin \Theta }{ \lambda } }{ (\frac{ \pi*a*\sin \Theta }{ \lambda })^2 }\]